John is contemplating on whether to go for a project that will give 100,000 each year for four years or a project that will give 75000 each year for six years. Advise him if the discount rate is 10%

1
Expert's answer
2022-02-24T08:15:35-0500

"PV= \frac{CF}{(1+r)^{n}}"

First Project

= "\frac{100000}{(1+0.1)^{4}}= \frac{100000}{1.4641}"

"= 68,301.35"


Second Project

"\frac{75000}{(1+0.1)^{6}}= \frac{75000}{1.771561}"

"=42,335.54"

I would advise him to under take the first project as it is more beneficial in terms of returns.